A chemist determines by measurements that 0.055 moles of oxygen gas participate in a chemical reaction.

Calculate the mass of oxygen gas that participates.

Round your answer to 2 significant digits.


1
Expert's answer
2021-08-26T01:08:16-0400

"n=0.055 \;moles \\

M(O_2) = 16+16 = 32 \;g/mol \\

m = n \times M = 0.055 \times 32 = 1.76 ≈ 1.8 \;g"

Answer: 1.8 g